### v1

# Rate-optimal and computationally efficient nonparametric estimation on the circle and the sphere

Abstract
We investigate the problem of density estimation on the unit circle and the unit sphere from a computational perspective. Our primary goal is to develop new density estimators that are both rate-optimal and computationally efficient for direct implementation. After establishing these estimators, we derive closed-form expressions for probability estimates over regions of the circle and the sphere. Then, the proposed theories are supported by extensive simulation studies. The considered settings naturally arise when analyzing phenomena on the Earth’s surface or in the sky (sphere), as well as directional or periodic phenomena (circle). The proposed approaches are broadly applicable, and we illustrate their usefulness through case studies in zoology, climatology, geophysics, and astronomy, which may be of independent interest. The methodologies developed here can be readily applied across a wide range of scientific domains.
Astronomy,
keywords:
[class=MSC]
keywords:
1 Introduction
Consider a random phenomenon expressed through some random variable $X$. The probability density function (PDF) $f=f_{X}$, or simply the density, of $X$ compresses all vital information for the random variable $X$, and consequently the target phenomenon. Expectations, event probabilities, and related quantitative insights are fully determined by the knowledge of $f$.
Of course, the density is unknown, and the most we can do is to “access” it from independent observations of it. In our context, we assume that $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ are independent realizations of $X$, which we must infer to assess the necessary information for $f$.
Nonparametric Statistics introduces estimators of $f$, defined as functions of the form $\widehat{f}_{n}:\mathcal{M}^{n}\times\mathcal{M}\to\mathbb{R}$, where $\mathcal{M}$ denotes the set of all possible values that the random variable $X$ can take. Given the random sample $X_{1},\dots,X_{n}$, we can construct $\widehat{f}_{n}(x)=\widehat{f}_{n}(X_{1},\dots,X_{n};x)$, $x\in\mathcal{M}$, which is referred to as the density estimator of $f$.
Nonparametric Statistics aims to provide such estimators without imposing restrictive—and often unrealistic—assumptions on the specific form of the unknown density. Instead, it seeks data-driven procedures that consistently estimate $f$ under mild conditions, such as continuity or smoothness/regularity. Typical constructions rely on kernels, wavelets, splines, and other sophisticated mathematical tools.
Historically, the earliest contributions are thanks to Rosenblatt [48], Parzen [42], and Breiman and Heller [4]. Foundational monographs include [50, 31], while the book [51] is now regarded as a standard modern reference. For a representative, but in no sense exhaustive, list of further developments in the field, see [1, 5, 6, 7, 11, 12, 17, 18, 19, 20, 21, 25, 26, 27, 28, 29, 30, 32, 33, 34, 37, 38, 39, 43, 44, 46, 47, 49].
In the present study we are interested in periodic random events and on random phenomena distributed on the Earth and Sky. The first can be modelled in terms of random variables distributed on the unit circle $\mathbb{S}^{1}=\{x\in\mathbb{R}^{2}:\|x\|=1\}$, while the second are expressed via random variables distributed on the unit sphere $\mathbb{S}^{2}=\{x\in\mathbb{R}^{3}:\|x\|=1\}$. Kernel density estimators (KDE) in that setting have a long history and applications in astrophysics; see for example [1], [2] and [22].
Recently, rate-optimal kernel density estimators in general metric spaces, covering the case of the spheres $\mathbb{S}^{d}$, $d\in\mathbb{N}$ have been established and explored in [10, 13, 14].
The main challenge here is that the theoretically optimal estimators, established in the aforementioned works, are expressed in terms of infinite series. For example for the case of the sphere $\mathbb{S}^{d}$, a “smooth” density $f$, can be estimated by the KDE
$\widehat{f}_{n,h}(x):=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{\infty}\frac{2\ell+1}{4\pi}g\Big(h\sqrt{\ell(\ell+1)}\Big)P_{\ell}(\langle x,X_{j}\rangle),\quad x\in\mathbb{S}^{2},$ (1.1)
for some appropriate function $g:\mathbb{R}\to\mathbb{R}$, and some bandwidth $h>0$. Note that $P_{\ell}$ in (1.1) are the well-known Legendre polynomials; for more see Section 2.3. This automatically means that the KDE cannot be used by any software in its stand form. Therefore, when we are in front of a data-study, we need to address the next:
Question: Can we propose new estimators, that are computationally-efficient and remain rate-optimal?
The answer to the above question is affirmative. The first problem we address in this study is to establish that there exists an index $N$ such that the finite-order estimator, obtained by summing from $\ell=0$ to $\ell=N$, retains rate optimality. Note that this corresponds to a frequency cutoff, which is a natural practise in Approximation Theory. The corresponding results are provided in Section 3.
Afterwards, with the established finite-order estimators in hand, we then turn our attention to probability estimation.
The new challenge here is that the numerical evaluation of the integral of a given density estimator is relatively slow and computationally intensive. Therefore, in order to make these probability estimators more suitable for use, we analytically integrate these estimators over a given region into a closed expression. Consequently, numerical computation is required only for evaluating these closed forms. Section 4 contains the derivation of the closed forms of these integrals and Figure 7 illustrates, at a glance, the resulting computational advantages.
In the sequel, we performed simulation studies to accompany these results. This can be found in Section 5 and showcases the accuracy of both the KDEs and probability estimators when compared to classic and mixture densities over the unit circle and sphere. This section also features evidence of the drastic time reduction in the evaluation of the closed forms of the probability estimators rather than the numerical integration method; see Section 5.3.
Section 6 presents four case studies in which our density estimators are applied to real-world data, with the aim not only of demonstrating the applicability of our theoretical results but also of disseminating the proposed methodology to the broader community.
Our first case study, drawn from zoology, concerns the movement directions of honeybees observed within a particular hive. As the data are naturally expressed as angles, our methods for the unit circle $\mathbb{S}^{1}$ are directly applicable.
Our second case study, arising from climatology, examines the annual precipitation patterns in Los Angeles, USA—a temporally periodic dataset. In such data, time can be interpreted as a periodic variable wrapped around the unit circle, making our methods on $\mathbb{S}^{1}$ particularly suitable. More generally, any periodic phenomenon can be treated by our methodologies presented here.
Our third case study, originating from geophysics, focuses on the spatial distribution of earthquake epicenters on the Earth’s surface. These locations are naturally represented as points on the unit sphere $\mathbb{S}^{2}$, making the use of our spherical estimators particularly appropriate for analyzing global seismic activity.
Our fourth case study, drawn from astrophysics, examines the distribution of bright stars in our galaxy. Each star is characterized by its galactic latitude and galactic longitude, providing data inherently defined on the unit sphere and thus ideally suited to our proposed estimators.
Finally, Section 8 (the Appendix) contains the proofs of all new theoretical results established in this study.
Notation.
Through all the document the sets $\mathbb{N}$, $\mathbb{N}_{0}$, denote the positive and non-negative integers while $\mathbb{R}$ and $\mathbb{C}$, the real and complex numbers respectively. By $g\in\mathcal{C}^{\tau}(\mathbb{R})$, we will mean that $g$ is a function in $\mathbb{R}$ with continuous derivatives up to order $\tau\in\mathbb{N}$. For any $s>0$, the number $\lfloor s\rfloor$ is the largest integer strictly smaller than $s$ and $\lceil s\rceil$ is the smallest integer strictly greater than $s$. We denote further $\mathcal{L}^{p}(\mathbb{S}^{d})=\{f:\mathbb{S}^{d}\to\mathbb{C}:\|f\|_{p}<\infty\}$, $p\in[1,\infty]$, where $\|f\|_{p}^{p}=\int_{\mathbb{S}^{d}}|f(x)|^{p}d\sigma_{\mathbb{S}^{d}}(x)$, for $p<\infty$ and $\|f\|_{\infty}=\text{essup}|f(x)|$, the Lebesgue spaces on $\mathbb{S}^{d}$. Finally $\langle x,y\rangle$ stands for the dot product of $x,y\in\mathbb{R}^{m}$, $m\in\mathbb{N}$.
2 Background
In this section, we collect the background material necessary for our study.
2.1 The underlying geometries
Firstly, let us formally introduce the settings of interest and their properties.
We start with $\mathbb{S}^{1}=\{x=(x_{1},x_{2})\in\mathbb{R}^{2}:\|x\|=1\}$ the unit circle. Recall the polar representation of each $x=(x_{1},x_{2})=(\cos\theta,\sin\theta)=:x(\theta)$, where $\theta$ is the unique angle in $(-\pi,\pi]$ given by
$\theta:=\begin{cases}\arccos x_{1},\quad x_{2}\geq 0\\ -\arccos x_{1},\quad x_{2}<0\end{cases}.$
The circle $\mathbb{S}^{1}$ is naturally suitable for the study of periodic functions. The circular distance, is the angle between the corresponding points and the origin;
$\rho_{\mathbb{S}^{1}}(x,y)=\arccos(\langle x,y\rangle)=\min\{|\theta-\phi|,2\pi-|\theta-\phi|\},\;x=x(\theta),y=y(\phi)\in\mathbb{S}^{1}.$
We denote by $d\sigma_{\mathbb{S}^{1}}(x)$ the length measure in $\mathbb{S}^{1}$. Then a measurable function $f:\mathbb{S}^{1}\to\mathbb{C}$ integrates as follows
$\int_{\mathbb{S}^{1}}f(x)d\sigma_{\mathbb{S}^{1}}(x)=\int_{-\pi}^{\pi}f(\cos\theta,\sin\theta)d\theta.$
Finally, the geometry is equipped with the (negative of the) Laplacian, which is given by $L_{\mathbb{S}^{1}}=-\frac{d^{2}}{d\theta^{2}}$.
We proceed to the sphere $\mathbb{S}^{2}=\{x=(x_{1},x_{2},x_{3})\in\mathbb{R}^{3}:\|x\|=1\}$. We present the necessary notions released from [52, Section 6] and [16, Section 1]. the spherical coordinates are given by
$x_{1}=\sin\theta\cos\phi,\quad x_{2}=\sin\theta\sin\phi,\quad x_{3}=\cos\theta,\quad(\theta,\phi)\in[0,\pi]\times(-\pi,\pi],$
and can be resolved as
$\theta=\arccos(x_{3}),\quad\text{and}\quad\phi=\begin{cases}\pi,\quad&\text{if }\;x_{1}=x_{2}=0,\\ \text{sign}(x_{2})\arccos\Big(\frac{x_{1}}{\sqrt{x_{1}^{2}+x_{2}^{2}}}\Big),\quad&\text{otherwise}\end{cases}.$
With a slight abuse of the notation we sometimes denote a point $x\in\mathbb{S}^{2}$ just by its spherical coordinates $x=(\theta,\phi)$.
We denote by $d\sigma_{\mathbb{S}^{2}}(x)$ the surface measure in $\mathbb{S}^{2}$. Then a measurable function $f:\mathbb{S}^{2}\to\mathbb{C}$ integrates as follows
$\int_{\mathbb{S}^{2}}f(x)d\sigma_{\mathbb{S}^{2}}(x)=\int_{-\pi}^{\pi}\int_{0}^{\pi}f(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)\sin\theta d\theta d\phi.$
The spherical distance is defined, as in $\mathbb{S}^{1}$, by
$\rho_{\mathbb{S}^{1}}(x,y)=\arccos(\langle x,y\rangle),\quad x,y\in\mathbb{S}^{2}.$
The (negative of the) spherical Laplacian in spherical coordinates is given by
$L_{\mathbb{S}^{2}}=-\frac{1}{\sin\theta}\frac{\partial}{\partial\theta}\Big(\sin\theta\frac{\partial}{\partial\theta}\Big)-\frac{1}{\sin^{2}\theta}\frac{\partial^{2}}{\partial^{2}\phi}.$
2.2 Legendre polynomials and spherical harmonics
Between the most valuable tools on the sphere, are the spherical harmonics and Legendre polynomials.
The Legendre polynomials $P_{\ell}$, $\ell\in\mathbb{N}_{0}$, are given by the simple Rodrigues formula
$P_{\ell}(u)=\frac{1}{2^{\ell}\ell!}\frac{d^{\ell}}{du^{\ell}}(u^{2}-1)^{\ell},\quad u\in[-1,1],\ell\in\mathbb{N}_{0}.$ (2.1)
We will further need the next precise expression of the Legendre polynomials:
$P_{\ell}(u)=\sum_{k=0}^{\ell}\binom{\ell}{k}\binom{\ell+k}{k}\frac{(-1)^{k}}{2^{k}}(1-u)^{k},\quad u\in[-1,1],\ell\in\mathbb{N}_{0},$ (2.2)
and the well-known upper bound
$|P_{\ell}(u)|\leq 1,\quad u\in[-1,1],\ell\in\mathbb{N}_{0}.$ (2.3)
Note further, that $P_{\ell}(1)=1$, for every $\ell\in\mathbb{N}_{0}$, $P_{0}\equiv 1$ and the set $\{P_{\ell}:\ell\in\mathbb{N}_{0}\}$ consist of an orthogonal basis for $L^{2}[-1,1]$, the space of square integrable functions on [1,1][-1,1].
The associated Legendre functions $P^{m}_{\ell}:\ell\in\mathbb{N}_{0},-\ell\leq m\leq m$, are defined by
$P_{\ell}^{m}(u):=\begin{cases}(-1)^{m}(1-u^{2})^{m/2}\frac{d^{m}}{du^{m}}P_{\ell}(u),\quad m\geq 0\\ (-1)^{m}\frac{(\ell+m)!}{(\ell-m)!}P_{\ell}^{-m}(u),\quad m<0\end{cases},\quad\ell\in\mathbb{N}_{0}.$ (2.4)
Of course $P_{\ell}^{0}=P_{\ell}$, for every $\ell\in\mathbb{N}_{0}$.
Finally, for every $\ell\in\mathbb{N}_{0}$ and $0\leq m\leq\ell$ the spherical harmonic $Y^{m}_{\ell}:[0,\pi]\times(-\pi,\pi]\to\mathbb{C}$ are given by
$Y_{\ell}^{m}(\theta,\phi)$ $=\sqrt{\frac{2\ell+1}{4\pi}}\sqrt{\frac{(\ell-m)!}{(\ell+m)!}}P_{\ell}^{m}(\cos\theta)e^{im\phi}=:N_{\ell,m}P_{\ell}^{m}(\cos\theta)e^{im\phi},$ (2.5)
and are expanded for $\ell\leq m<0$ by the next conjugation;
$Y^{m}_{\ell}(\theta,\phi)=(-1)^{m}\overline{Y^{-m}_{\ell}(\theta,\phi)},\quad\ell\in\mathbb{N}_{0},\;-\ell\leq m\leq 0.$ (2.6)
The most important property of spherical harmonics is that their union $\{Y^{m}_{\ell}:\ell\in\mathbb{N}_{0},-\ell\leq m\leq\ell\}$ consists of an orthonormal basis of $(\mathcal{L}^{2}(\mathbb{S}^{2}),d\sigma_{\mathbb{S}^{2}})$. Moreover, they are connected with Legendre polynomials under the fundamental additive formula:
$\frac{2\ell+1}{4\pi}P_{\ell}(\langle x,y\rangle)=\sum_{m=-\ell}^{\ell}Y^{m}_{\ell}(\theta_{x},\phi_{x})\overline{Y_{\ell}^{m}(\theta_{y},\phi_{y})},\quad x,y\in\mathbb{S}^{2},\;\ell\in\mathbb{N}_{0}.$ (2.7)
2.3 Density estimation on the circle and sphere
Let us now summarise some key results from [10],[13], and [14], in which density estimation has been studied on metric spaces $\mathcal{M}$ associated with operators $L$ under mild assumptions. In our case, the spaces $\mathbb{S}^{d}$, for $d=1,2$, with corresponding operators, the negative Laplacians $L_{\mathbb{S}^{d}}$, given in Section 2.1, obey these assumptions and therefore the results from those studies hold automatically. We present those results in the case of $\mathbb{S}^{d}$, for $d=1,2$.
Let $g:\mathbb{R}\to\mathbb{R}$ be a measurable function such that $g(0)=1$ with a “fast” decay (this is specified in Theorem 2.1). Such a function is referred as a symbol and will generate a KDE, as in the case of kernel functions on $\mathbb{R}^{d}$. Let $h>0$ a quantity that extends the notion of a bandwidth.
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$.
Let $d=1$. The KDE $\widehat{f}_{n,h}$ takes the following form:
$\widehat{f}_{n,h}(x)=\frac{1}{2\pi n}\sum_{j=1}^{n}\left(1+2\sum_{\ell=1}^{\infty}g(h\ell)\cos(\ell(\theta-\theta_{j}))\right),\quad x=x(\theta),\;\theta\in(-\pi,\pi],$ (2.8)
where $\theta_{j}$ the polar angles of $X_{j}$, $j=1,\dots,n$ and $\theta$ the polar angle of the arbitrary point $x\in\mathbb{S}^{1}$.
Let $d=2$. The KDE $\widehat{f}_{n,h}$ takes the following form:
$\widehat{f}_{n,h}(x)=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{\infty}\frac{2\ell+1}{4\pi}g\left(h\sqrt{(\ell(\ell+1))}\right)P_{\ell}(\langle x,X_{j}\rangle),\quad x\in\mathbb{S}^{2}.$ (2.9)
As always in Non-parametric statistics, we assume that the unknown density, lies on a regularity space. Let $s>0$. We say that a function $f$ belongs to the Hölder space of regularity order $s$, $\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$, when
$\|f\|_{\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})}=\|f\|_{\dot{\mathcal{H}}^{s}}=\sup_{x\neq y}\frac{|L_{\mathbb{S}^{d}}^{k/2}f(x)-L_{\mathbb{S}^{d}}^{k/2}(y)|}{\rho_{\mathbb{S}^{d}}(x,y)^{s-k}}<\infty,$
where $k$ the smallest integer strictly less than $s$.
Observe that $L_{\mathbb{S}^{d}}$ are second order differential operators, which indicates their use for defining regularity spaces and the half in the exponent. For the fractional powers of the operators, we refer the reader to [45, p. 263]. For more details on regularity spaces on the sphere and more general spaces, we refer to [15, 24, 23, 35, 36]
Finally, the simplest measure of the accuracy of an estimator is the mean squared error:
$\text{MSE}($ $\widehat{f}_{n,h})(x):=\mathbb{E}_{f}[(\widehat{f}_{n,h}(x)-f(x))^{2}]$
$=\int_{\mathbb{S}^{d}}\dots\int_{\mathbb{S}^{d}}(\widehat{f}_{n,h}(x;x_{1},\dots,x_{n})-f(x))^{2}f(x_{1})\dots f(x_{n})\times$
$\hskip 85.35826pt\times d\sigma_{\mathbb{S}^{d}}(x_{1})\dots d\sigma_{\mathbb{S}^{d}}(x_{n}).$ (2.10)
By Theorem 3.1 in [14] we infer:
Theorem 2.1.
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g:\mathbb{R}\to\mathbb{R}$ be an even measurable function such that $g\in\mathcal{C}^{\tau}(\mathbb{R})$, for some $\tau>d+s$, $g(0)=1$, $g^{(\nu)}(0)=0$, for $1\leq\nu\leq\tau$, and for some $r>\tau+d$ and $C_{\tau}>0$,
$|g^{(\nu)}(\lambda)|\leq C_{\tau}(1+|\lambda|)^{-r},\quad\text{for every}\;\lambda\in\mathbb{R},\;\text{and}\;0\leq\nu\leq\tau.$
Then the corresponding KDEs $\widehat{f}_{n,h}$ as in (2.8), for $d=1$, and (2.9) for $d=2$ satisfy,
$\sup_{x\in\mathbb{S}^{d}}\textnormal{MSE}(\widehat{f}_{n,h})(x)\leq cC(f)n^{-2s/(2s+d)},$ (2.11)
where $c>0$ depends only on $\tau,\,s,\,C_{\tau}$ and $d$, while $C(f)=\max\{\|f\|_{\infty},\|f\|^{2}_{\dot{\mathcal{H}}^{s}}\}$.
The reader is referred to [10, 13] for more general results in terms of errors and function spaces and to [1, Theorem 11] for the lower bounds and thus the optimality of the rate above.
3 Rate-optimal and computationally efficient estimators
Whilst Theorem 2.1 shows that the KDEs in (2.8) and (2.9) are theoretically suitable estimators, the implementation of these KDEs comes with a computational challenge; each expression contains an infinite sum over $\ell$. To evaluate these estimators in practice, we must replace the infinite series by a finite-order version, keeping only terms up to some cutoff index $N$. Determining this index $N$ in such a way that the estimator remains rate-optimal motivates us to establish the next theorem, whose proof, as all proofs of our study, can be found in the Appendix:
Theorem 3.1.
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid copies of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$. Let also $h=h_{n}=n^{-1/(2s+d)}$ and $g:\mathbb{R}\to\mathbb{R}$ be an even, measurable function such that $g\in\mathcal{C}^{\tau}(\mathbb{R})$, for some $\tau>d+s$, $g(0)=1$, $g^{(\nu)}(0)=0$, for $1\leq\nu\leq\tau$, and for some $r>\tau+d$ and $C_{\tau}>0$,
$|g^{(\nu)}(\lambda)|\leq C_{\tau}(1+|\lambda|)^{-r},\quad\text{for every}\;\lambda\in\mathbb{R},\;\text{and}\;0\leq\nu\leq\tau.$ (3.1)
We denote by
$N_{s}:=\left\lfloor\left(\dfrac{n^{(s+r)/(2s+d)}}{d\pi(r-d)}\right)^{1/(r-d)}\right\rfloor+1.$ (3.2)
We introduce the following —finite-order— estimators:
$\bullet$ When $d=1$
$\widehat{f}_{s}(x):=\frac{1}{2\pi n}\sum_{j=1}^{n}\Big(1+2\sum_{\ell=1}^{N_{s}}g(h\ell)\cos(\ell(\theta-\theta_{j}))\Big),\quad x=x(\theta)\in\mathbb{S}^{1},$ (3.3)
where $\theta,\theta_{j}\in(-\pi,\pi]$ represent the polar angles of the arbitrary $x\in\mathbb{S}^{1}$ and $X_{j}$, $j=1,\dots,n$, respectively.
$\bullet$ When $d=2$
$\widehat{f}_{s}(x):=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{N_{s}}\frac{2\ell+1}{4\pi}g\big(h\sqrt{\ell(\ell+1)}\big)P_{\ell}(\langle x,X_{j}\rangle),\quad x\in\mathbb{S}^{2}.$ (3.4)
Then the estimator $\widehat{f}_{s}$ as in (3.3) for $d=1$ and in (3.4) for $d=2$ enjoys
$\sup_{x\in\mathbb{S}^{d}}\textnormal{MSE}(\widehat{f}_{s})(x)\leq C^{\prime}n^{-2s/(2s+d)},$ (3.5)
where $C^{\prime}$ depends on $C_{\tau}$ as well as $c$ and $C(f)$ as in Theorem 2.1.
We close the present section by listing the next remarks.
Remark 3.1.
(a) Our finite-order KDES in (3.3) and (3.4) of Theorem 3.1, are computationally efficient, once we fix the appropriate symbol $g$. Let $d\in\{1,2\}$ and $s>0$ and set $r:=2d+\lceil s\rceil+1$. Then the symbol
$g_{r}(\lambda):=\frac{1}{1+|\lambda|^{r}},\quad\lambda\in\mathbb{R},$ (3.6)
satisfies the assumptions of Theorem 3.1; see also [13, 14]. This means that our new optimal estimators are ready-to-go for any data analysis. For simulations and applications in the current study we refer the reader to the Sections 5 and 6.
(b) The finite estimators in Theorem 3.1, indeed integrate to the unit; $\int_{\mathbb{S}^{d}}\widehat{f}_{s}(x)d\sigma_{\mathbb{S}^{d}}(x)=1$, as one would expect. For the case of the circle $\mathbb{S}^{1}$, we can confirm this just by integrating (3.3) and using the properties of the trigonometric functions. For the case of the sphere $\mathbb{S}^{2}$, we just have to use the Funk–Hecke formula; [16, Theorem 1.2.9] to obtain for every $y\in\mathbb{S}^{2}$,
$\int_{\mathbb{S}^{2}}P_{\ell}(\langle x,y\rangle)d\sigma_{\mathbb{S}^{2}}(x)=2\pi\int_{-1}^{1}P_{\ell}(u)du=\begin{cases}4\pi,\quad\ell=0\\ 0,\quad\ell\neq 0\end{cases},$
with the last holding because $P_{0}\equiv 1$ and thanks to the orthogonality of the Legendre polynomials.
(c) The KDEs obtained here, in the language of Harmonic Analysis, are commonly referred as band-limited. This is because they are obtained by Fourier expansions, using a finite part of the spectrum of the associated operators (Laplacians). This is a very standard direction in Analysis and Approximation Theory and finds applications in a wide range of areas.
4 Probability estimation
A problem of fundamental interest is the estimation of the probability of an event occurring within a certain region of a space. Our kernel density estimators can be used for such an application, not only as plug-in estimators but also as a rate optimal estimation for such a probability on the circle or sphere. We introduce the following probability estimators for every measurable set $A\subset\mathbb{S}^{d}$, $d=1,2$ and we prove that these can indeed well-estimate the probability of any measurable subset $A$.
Proposition 4.1.
Consider a random variable $X$ distributed on $\mathbb{S}^{d}$, $d\in\{1,2\}$, with unknown density $f$ and $X_{1},\dots,X_{n}$, $n\in\mathbb{N}$ iid samples of $X$. Assume that $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{d})$ for some $s>0$ and denote by $\widehat{f}_{s}$ the finite KDE as in Theorem 3.1. Let finally $A\subset\mathbb{S}^{d}$ be a measurable set. Then the following probability estimator
$\widehat{\mathbb{P}}(A):=\widehat{\mathbb{P}}(A,X_{1},\dots,X_{n}):=\int_{A}\widehat{f}_{s}(x)d\sigma_{\mathbb{S}^{d}}(x),$ (4.1)
satisfies
$\mathbb{E}_{f}[(\widehat{\mathbb{P}}(A)-\mathbb{P}(A))^{2}]\leq c^{\prime}n^{-2s/(2s+d)},$
where $c^{\prime}>0$ depends on the measure of $\sigma_{\mathbb{S}^{d}}(A)$ as well as $C^{\prime}$ from Theorem 3.1.
With Proposition 4.1 in hand we can integrate our new estimators and obtain probabilities of certain regions of $\mathbb{S}^{d}$. We start with the case of $d=1$. Let $\pi<\vartheta_{1}<\vartheta_{2}\leq\pi$ and we introduce the following angular domain
$A[\vartheta_{1},\vartheta_{2}]=\{x=x(\theta)\in\mathbb{S}^{1}:\;\vartheta_{1}\leq\theta\leq\vartheta_{2}\}.$ (4.2)
We apply Proposition 4.1 to get the following:
Proposition 4.2.
Let $X$ a random variable distributed on $\mathbb{S}^{1}$, with density $f\in\mathcal{L}^{\infty}(\mathbb{S}^{1})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{1})$ for some $s>0$ and $X_{1},\dots,X_{n}$ iid samples of $X$. For any function $g:\mathbb{R}\to\mathbb{R}$ as in Theorem 3.1 and $h=n^{-1/(2s+1)}$. Then
$\widehat{\mathbb{P}}(A[\vartheta_{1},\vartheta_{2}])$ $=\frac{\vartheta_{2}-\vartheta_{1}}{2\pi}$ (4.3)
$+\frac{1}{\pi n}\sum_{j=1}^{n}\sum_{\ell=1}^{N_{s}}\frac{g(h\ell)}{\ell}\left(\sin(\ell(\vartheta_{2}-\theta_{j}))-\sin(\ell(\vartheta_{1}-\theta_{j}))\right),$
where $\theta_{j}$ the polar angles of $X_{j}$, $j=1,\dots,n$.
The integration of the finite estimators (3.3) on the circle, is straightforward; see the Appendix. In contrast, our estimators (3.4) for $\mathbb{S}^{2}$ involve the Legendre polynomials and their integration turns out to be challenging. Numerical methods end up taking a very long time to compute integrals of the form
$\int_{A}P_{\ell}(\langle x,x_{j}\rangle)d\sigma_{\mathbb{S}^{2}}(x),\quad A\subset\mathbb{S}^{2},$
despite how simple form the domain $A$ has, and therefore the software suffers providing the final estimated probabilities.
This motivates the direct computation of the integral of the finite estimator (3.4), as we will do in the next theorem. We will see that this closed form is also particularly suitable for vectorisation when being evaluated, further decreasing run-time, especially when combined with parallel programming. The anti-derivatives of the form of the Legendre polynomials in (3.4) can be written in a closed form but special functions will be needed in this expression. These special functions are well-known and are found natively in software like Mathematica or are easily available through standard packages, such as the scipy package for Python.
We will particularly need the so-called incomplete Beta function. For every $a,b>0$ the incomplete Beta function is defined by
$B(u;a,b):=\int_{0}^{u}t^{a-1}(1-t)^{b-1}dt,\quad\text{for every }u\in[0,1].$ (4.4)
Note that for $u=1$, its value is equal to the standard Beta function. Furthermore, if $a,b\in\mathbb{N}$, then the integrand function in (4.4) is a polynomial, so the incomplete Beta function, its anti-derivative, is just a polynomial too. For the better presentation of the study we introduce the next kernels: For every $\ell\in\mathbb{N}$, $1\leq m\leq k\leq\ell$ and $\gamma_{1},\gamma_{2}\in[-1,1]$, we denote by
$\mathcal{B}_{m,k}(\gamma_{1},\gamma_{2}):=B\Big(\frac{1+\gamma_{2}}{2};\frac{m}{2}+1,k-\frac{m}{2}+1\Big)-B\Big(\frac{1+\gamma_{1}}{2};\frac{m}{2}+1,k-\frac{m}{2}+1\Big).$ (4.5)
As in the case of the circle we work with angular domains; namely denote $\vartheta=(\vartheta_{1},\vartheta_{2}),\;\varphi=(\varphi_{1},\varphi_{2})$, where $0\leq\vartheta_{1}<\vartheta_{2}\leq\pi$ and $\pi\leq\varphi_{1}<\varphi_{2}\leq\pi$. We set
$\mathcal{A}_{\vartheta,\varphi}:=\{x=x(\theta,\phi)\in\mathbb{S}^{2}:(\theta,\phi)\in[\vartheta_{1},\vartheta_{2}]\times[\varphi_{1},\varphi_{2}]\}.$
Theorem 4.1.
Let $X$ a random variable distributed on $\mathbb{S}^{2}$, with density $f\in\mathcal{L}^{\infty}(\mathbb{S}^{2})\cap\dot{\mathcal{H}}^{s}(\mathbb{S}^{2})$ for some $s>0$ and $X_{1},\dots,X_{n}$ iid samples of it. For any function $g:\mathbb{R}\to\mathbb{R}$ as in Theorem 3.1 and $h=n^{-1/(2s+2)}$. Then for every $\vartheta=(\vartheta_{1},\vartheta_{2}),\;\varphi=(\varphi_{1},\varphi_{2})$, with $0\leq\vartheta_{1}<\vartheta_{2}\leq\pi$ and $\pi\leq\varphi_{1}<\varphi_{2}\leq\pi$, we have that
$\widehat{\mathbb{P}}(\mathcal{A}_{\vartheta,\varphi})=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{N_{s}}g\Big(h\sqrt{\ell(\ell+1)}\Big)\Big(I_{\ell}^{0}(j)+2\sum_{m=1}^{\ell}I_{\ell}^{m}(j)\Big),$ (4.6)
where:
for every $\ell\in\mathbb{N}_{0}$ and $j=1,\dots,n$,
$I_{\ell}^{0}(j)=N_{\ell,0}^{2}(\varphi_{2}-\varphi_{1})P_{\ell}(\cos\theta_{j})\sum_{k=0}^{\ell}$ ${\binom{\ell}{k}\binom{\ell+k}{k}}\left(\frac{-1}{2}\right)^{k}\cdot\frac{1}{k+1}$
$\left((1-\cos\vartheta_{2})^{k+1}-(1-\cos\vartheta_{1})^{k+1}\right);$
for every $\ell\in\mathbb{N}$, $1\leq m\leq\ell$ and $j=1,\dots,n$,
$I^{m}_{\ell}(j)$ $=\frac{1}{m}N_{\ell,m}^{2}P^{m}_{\ell}(\cos\theta_{j})\Big(\sin\big(m(\varphi_{2}-\phi_{j})\big)-\sin\big(m(\varphi_{1}-\phi_{j})\big)\Big)\times$
$\hskip 17.07182pt\times\sum_{k=m}^{\ell}{\binom{\ell}{k}\binom{\ell+k}{k}}\frac{2(-1)^{k}k!}{(k-m)!}\mathcal{B}_{k,m}(\cos\vartheta_{2},\cos\vartheta_{1})$
for $N_{\ell,m}$ as in (2.5), $\mathcal{B}_{k,m}$ as in (4.5) and where $(\theta_{j},\phi_{j})$ are the spherical coordinates of the observations $X_{j}$, $j=1,\dots,n$.
The proof of the above theorem, relies on the next result, which is of potential independent interest, beyond the field of Statistics.
Lemma 4.2.
Let $1\leq\gamma_{1}<\gamma_{2}\leq 1$, $\ell\in\mathbb{N}$ and $1\leq m\leq\ell$. Then the associated Legendre functions $P_{\ell}^{m}$ integrate in the closed interval $\gamma_{1},\gamma_{2}]$ as
$\int_{\gamma_{1}}^{\gamma_{2}}P_{\ell}^{m}(u)du$ $=\sum_{k=m}^{\ell}\binom{\ell}{k}\binom{\ell+k}{k}\frac{2(-1)^{k}k!}{(k-m)!}\mathcal{B}_{k,m}(\gamma_{1},\gamma_{2})$
where $\mathcal{B}_{k,m}$ the kernels in (4.5) involving the incomplete Beta function (4.4).
5 Simulations
In this section, we reinforce with simulations the theoretical finds in Sections 3 and 4. We simulate from classic distributions on the sphere and circle and present our estimation of the densities and probabilities; that is the implementation of Theorem 3.1, Proposition 4.2 and Theorem 4.1. In particular, we first sample from the uniform distributions on $\mathbb{S}^{d}$ and then the well-known von Mises-Fisher distribution.
Our methodology for this section is as follows: we select a random variable on $\mathbb{S}^{1}$ or $\mathbb{S}^{2}$ with a density $f$ and simulate $n=1000$ points from its distribution. We then construct finite-order kernel density estimators $\widehat{f}_{s}$, as in Theorem 3.1, taking $s=0.5$, 11, and $2$ with $r=2d+\lceil s\rceil+1$ and symbol $g_{r}$, as in Remark 3.1. We then plot the original distribution $f$ alongside the simulated sample of $n=1000$ points and the KDEs $\widehat{f}_{0.5}$, $\widehat{f}_{1}$, and $\widehat{f}_{2}$.
In the case of $\mathbb{S}^{1}$, the plot is given on an annulus (with an inner radius of $0.8$ and outer radius of $1.2$) with the colour representing the value of the function at a given angle, achieved with the ColorFunction option in Mathematica’s PolarPlot. The points are also given on the same annulus with the angle of each point in polar coordinates being equal to the angle of the observation but with the radius given as a random number between $0.8$ and $1.2$ in order to better spread out the points for visual clarity.
In the case of $\mathbb{S}^{2}$, the plot is given on a 3D rendering of a sphere, viewed from the point (1.3,2.4,2)(1.3,-2.4,2), the default in Mathematica’s SphericalPlot3D. The spheres are coloured according to their corresponding function (the PDF or KDE). This colour plot was achieved by evaluating each function over a Table in terms of $\theta$ and $\phi$. In these plots we took the step size $\pi/32$ in each coordinate giving a table of $65\times 33$ points which were entered into ListDensityPlot to get an Image which was used as the Texture in the PlotStyle option of SphericalPlot3D. The $n=1000$ points are given on the same sphere but with no added colour.
For each simulated KDE, we estimated the probability over several regions using the methods established in Proposition 4.2 and Theorem 4.1 for $\mathbb{S}^{1}$ and $\mathbb{S}^{2}$, respectively. Precisely, the Tables 1-5 contain: these estimated probabilities compared to the observed frequency in the given region as well as the true (theoretical) probability, obtained by integrating the known PDF over the same region.
Finally, alongside the probabilities in these tables, we also include an indicator of the performance of each KDE: the estimated mean integrated square error (MISE). The MISE is a global criterion for the fit of an estimator $\widehat{f}$ to a density $f$ and is given by
$\text{MISE}(\widehat{f})=\mathbb{E}\left[\int_{\mathbb{S}^{d}}(\widehat{f}(x)-f(x))^{2}d\sigma_{\mathbb{S}^{d}}(x)\right],\quad d=1,2.$ (5.1)
Note also that, by the Fubini-Tonelli theorem and Theorem 3.1, we have when $f\in\mathcal{L}^{\infty}(\mathbb{S}^{d})\cap\dot{\mathcal{H}}(\mathbb{S}^{d})$, for some $s>0$ and $\widehat{f}_{s}$ as in Theorem 3.1,
$\text{MISE}(\widehat{f}_{s})\leq\sigma_{\mathbb{S}^{d}}(\mathbb{S}^{d})C^{\prime}n^{-2s/(2s+d)},\quad d=1,2.$ (5.2)
In order to estimate the MISE, we simulated $n=1000$ points, $30$ times, computed the MISE in each instance and took the mean of this sample.
The form of the results of Proposition 4.2 and Theorem 4.1 are suitable for array programming. This means that for a given region of $\mathbb{S}^{1}$ or $\mathbb{S}^{2}$ to integrate over, one can apply functions such as the associated Legendre polynomials $P_{m}^{\ell}$ over the entire dataset for all necessary values of $m$ and $\ell$ simultaneously. This drastically reduces computation time, especially when compared to numerical integration (see Section 5.3). Finally, a computational tip is that there is no need to calculate the inclination angles $\theta_{j}$ as this only appears as the argument of $\cos$ and by (2.1) $\cos\theta_{j}=X_{j3}$, the third coordinate of the $j^{\text{th}}$ observation.
5.1 The uniform distribution
We begin with simulations of the uniform distribution on the unit sphere. In order to simulate from this distribution on $\mathbb{S}^{d}$, the usual approach, from [41] is given in Algorithm 1.
Algorithm 1 Simulation of uniformly distributed points on $\mathbb{S}^{d}$
1:Dimension of the sphere: $d\in\mathbb{N}$
2:Sample size: $n\in\mathbb{N}$
3:Create $(n\times(d+1))$ matrix $\mathbf{A}=(a_{ij})$
[…]
5:while $i\leq n$ do
6:  Independently sample $Z_{1},\dots,Z_{d+1}\sim N(0,1)$
7:  $j\leftarrow 1$
8:  while $j\leq d+1$ do
9:   $a_{ij}\leftarrow Z_{j}/\sqrt{Z_{1}+\dots+Z_{d+1}}$
10:   $j\leftarrow j+1$
11:  end while
12:  $i\leftarrow i+1$
13:end while
14:return matrix $\mathbf{A}$, each row vector of which will follow a uniform distribution on $\mathbb{S}^{d}$
We plot and evaluate our new KDEs in Figure 1 and Table 1.
Figure 1: (Left to right) The plot of the uniform distribution over $\mathbb{S}^{2}$; a sample of $n=1000$ points from such a distribution; KDEs constructed using said data for $s=0.5$, $s=1$ and $s=2$, respectively.
The plots indicate that the estimators have captured the simulated data.
The KDE estimated probabilities (taken over halves and quarters of $\mathbb{S}^{2}$) shown in Table 1 are all comparable with both the observed frequencies and true probabilities, confirming the performance of the estimators introduced in the paper.
Note finally that a preference to larger smoothness levels, say $s\in[1,2]$, seems better for such a uniform data, because it can avoid overfitting the data and ends up to very low MISE averages.
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
$N_{s}=19$ $N_{s}=8$ $N_{s}=4$
Half 1 $0,\pi]\times(-\pi,0]$ $0.5100$ $0.5080$ $0.5077$ $0.518$ $0.5$
Half 2 $0,\pi]\times[0,\pi]$ $0.4900$ $0.4920$ $0.4923$ $0.482$ $0.5$
Quarter 1 $0,\pi/2]\times(-\pi,0]$ $0.2454$ $0.2439$ $0.2422$ $0.246$ $0.25$
Quarter 2 $0,\pi/2]\times[0,\pi]$ $0.2471$ $0.2495$ $0.2464$ $0.243$ $0.25$
Quarter 3 $\pi/2,\pi]\times(-\pi,0]$ $0.2646$ $0.2641$ $0.2655$ $0.272$ $0.25$
Quarter 4 $\pi/2,\pi]\times[0,\pi]$ $0.2429$ $0.2425$ $0.2459$ $0.239$ $0.25$
MISE $0.00643$ $0.00204$ $0.00062$ - -
Table 1: Estimated probabilities from a uniform distribution on $\mathbb{S}^{2}$ for $n=1000$ points and $s=0.5$, $s=1$, $s=2$, as well as the corresponding estimated MISE.
Analogously, Figure 2 and Table 2 contain the plots of the KDEs and the associated probabilities with the estimated MISE on $\mathbb{S}^{1}$. Again the performance is satisfactory and $s\in[1,2]$ seems enough to avoid overfitting and offering low MISEs.
Figure 2: (Left to right) The plot of the uniform distribution over $\mathbb{S}^{1}$; a sample of $n=1000$ points from such a distribution; KDEs constructed using said data for $s=0.5$, $s=1$, and $s=2$, respectively.
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
$N_{s}=85$ $N_{s}=17$ $N_{s}=6$
Half 1 $(-\pi,0]$ $0.5124$ $0.5153$ $0.5130$ $0.515$ $0.5$
Half 2 $0,\pi]$ $0.4876$ $0.4847$ $0.4870$ $0.485$ $0.5$
Quarter 1 $(-\pi,-\pi/2]$ $0.2653$ $0.2676$ $0.2639$ $0.266$ $0.25$
Quarter 2 $\pi/2,0]$ $0.2471$ $0.2477$ $0.2491$ $0.249$ $0.25$
Quarter 3 $0,\pi/2]$ $0.2372$ $0.2394$ $0.2387$ $0.233$ $0.25$
Quarter 4 $\pi/2,\pi]$ $0.2504$ $0.2453$ $0.2483$ $0.252$ $0.25$
MISE $0.00799$ $0.00248$ $0.00092$ - -
Table 2: Estimated probabilities from a uniform distribution on $\mathbb{S}^{1}$ for $n=1000$ points and $s=0.5$, $s=1$, $s=2$, and corresponding estimated MISE.
5.2 The von Mises-Fisher’s distribution
We shall now move to estimating samples from the von Mises-Fisher (vMF) distribution; the spherical analogue of the normal distribution.
Recall that the vMF distribution depends on two parameters: the mean direction $\mu\in\mathbb{S}^{d}$ and the shape $\kappa>0$. Its density its given by
$f(x)=c_{d,\kappa}\exp\big(\kappa\langle x,\mu\rangle\big),\quad x\in\mathbb{S}^{d},$ (5.3)
where $c_{d,\kappa}$ is the normalisation constant. From [40], we have that for $d=1$, this constant does not have a closed form and involves the modified Bessel function of the first kind: $c_{1,\kappa}=1/(2\pi I_{0}(\kappa))$. However, for $d=2$, one readily finds that the constant is $c_{2,\kappa}=\kappa/(4\pi\sinh(\kappa))$.
The vMF density $f$ has a single peak at $\mu$ and the decay as $x$ departs from $\mu$ depends only on the distance between $x$ and $\mu$, see (2.1), and the value of $\kappa$. Indeed, $\kappa$ plays the role of the reciprocal of the variance in a univariate Gaussian distribution with larger values of $\kappa$ corresponding to a greater concentration of the density about $\mu$. As $\kappa\to 0$, the vMF distribution approaches the spherical uniform distribution.
We now briefly present the methods we used to simulate from the vMF distribution. In the following, we assume the vMF distribution has a shape parameter $\kappa>0$ and mean direction $\mu=(1,0)$, in the $\mathbb{S}^{1}$ case, or $\mu=(0,0,1)$, in the $\mathbb{S}^{2}$ case. For other mean directions, the data need only be appropriately rotated after sampling.
The algorithm, presented in [3], for sampling over the vMF distribution on $\mathbb{S}^{1}$, after fixing $\kappa>0$, is as given in Algorithm 2.
Algorithm 2 Simulation of von Mises-Fisher distributed points on $\mathbb{S}^{1}$
1:Mean direction vector: $\mu=(\mu_{1},\mu_{2})\in\mathbb{S}^{1}$
2:Concentration parameter: $\kappa>0$
3:Sample size: $n\in\mathbb{N}$
4:Create $(n\times 2)$ matrix $\mathbf{M}=(m_{ij})$
[…]
9:while $i\leq n$ do
10:  $p\leftarrow 0$
11:  $q\leftarrow 0$
12:  while $p\leq 0$ and $q\leq 0$ do
13:   Independently sample $U_{1},U_{2},U_{3}\simUnif$$[0,1]$
14:   $z\leftarrow\cos(\pi U_{1})$
15:   $f\leftarrow(1+rz)/(r+z)$
16:   $c\leftarrow(r-f)\kappa$
17:   $p\leftarrow c(2-c)U_{2}$
18:   $q\leftarrow\ln(c/U_{2})+1-c$
19:  end while
20:  $\vartheta\leftarrowsign$$(U_{3}-1/2)\arccos f$
21:  $m_{i1}\leftarrow\cos\vartheta$
22:  $m_{i2}\leftarrow\sin\vartheta$ $\triangleright$ The rows of $\mathbf{M}$ will be vMF with mean $(1,0)$
23:  $i\leftarrow i+1$
24:end while
25:if $\mu_{2}\geq 0$ then
26:  $\phi\leftarrow\arccos\mu_{1}$
[…]
28:end if
[…]
In order to simulate from the von Mises-Fisher distribution on $\mathbb{S}^{2}$, we follow the process given in [53], which is given in Algorithm 3.
Algorithm 3 Simulation of von Mises-Fisher distributed points on $\mathbb{S}^{2}$
1:Mean direction vector: $\mu=(\mu_{1},\mu_{2},\mu_{3})\in\mathbb{S}^{2}$
2:Concentration parameter: $\kappa>0$
3:Sample size: $n\in\mathbb{N}$
4:Create $(n\times 3)$ matrix $\mathbf{M}=(m_{ij})$
[…]
6:while $i\leq n$ do
7:  Sample $\varphi\simUnif$$(-\pi,\pi]$
8:  Sample $U\simUnif$$[0,1]$
9:  $\vartheta\leftarrow\arccos(1+\frac{1}{\kappa}\ln(U+(1-U)e^{-2\kappa}))$
10:  $m_{i,1}\leftarrow\cos\varphi\sin\vartheta$
11:  $m_{i,2}\leftarrow\sin\varphi\sin\vartheta$
12:  $m_{i,3}\leftarrow\cos\vartheta$ $\triangleright$ The rows of $\mathbf{M}$ will be vMF with mean $(0,0,1)$
13:  $i\leftarrow i+1$
14:end while
15:if $\mu\neq(0,0,-1)$ then
16:  $\mathbf{R}\leftarrow\begin{pmatrix}1-\frac{\mu_{1}^{2}}{1+\mu_{3}}&-\frac{\mu_{1}\mu_{2}}{1+\mu_{3}}&-\mu_{1}\\ -\frac{\mu_{1}\mu_{2}}{1+\mu_{3}}&1-\frac{\mu_{2}^{2}}{1+\mu_{3}}&-\mu_{2}\\ \mu_{1}&\mu_{2}&\mu_{3}\end{pmatrix} \triangleright$ Make a rotation matrix
[…]
18:end if
[…]
Implementation
In Figure 3 we plot the $\kappa=1$, $\mu=(0,0,1)$ vMF distribution alongside the $n=1000$ sample and three KDEs for $s=0.5$, 11, 22. Table 3 contains the observed frequencies and true probabilities over four regions centred on the mean direction, alongside the MISE corresponding to each value of $s$.
Figure 3: (Left to right) The plot of the vMF distribution over $\mathbb{S}^{2}$ with $\kappa=1$ and $\mu=(0,0,1)$; a sample of $n=1000$ points from such a distribution; KDEs constructed using said data for $s=0.5$, $s=1$ and $s=2$, respectively.
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
$N_{s}=19$ $N_{s}=8$ $N_{s}=4$
$0,\pi/2]\times(-\pi,\pi]$ 0.7398 0.7414 0.7435 0.736 0.7311
$0,\pi/3]\times(-\pi,\pi]$ 0.4578 0.4573 0.4576 0.45 0.4551
$0,\pi/4]\times(-\pi,\pi]$ 0.2913 0.2964 0.2925 0.289 0.2936
$0,\pi/5]\times(-\pi,\pi]$ 0.2028 0.2058 0.1991 0.199 0.2011
MISE 0.00672 $0.00203$ $0.00063$ - -
Table 3: Estimated probabilities from a vMF ($\kappa=1$, $\mu=(0,0,1)$) distribution for $n=1000$ points and $s=0.5$, $s=1$, $s=2$.
We can see that all values of $s$ gave very accurate estimations of the probabilities over each region. The plots and MISE both indicate that $s=2$ gives the best fit out of the values we considered. See Figure 4 for the MISE as $s$ ranges over $\{0.1,0.2,\dots,10\}$ where we see an approximate local minimum at $s=2.3$. We will see in the next examples that smaller values of $s$ will be more appropriate for densities with more local behaviour, and we will later use this conclusion in Section 6.3.
Figure 4: The estimated MISE for vMF simulated points with $\kappa=1$ and $s\in\{0.1,0.2,\dots,10\}$, showing a clear minimum around $s=2$.
Let us now consider von Mises-Fisher mixture densities. More precisely, consider some $m\in\mathbb{N}$ and positive weights $w_{1},\dots,w_{m}$ such that $\sum_{j}w_{j}=1$. Then for mean directions $\mu_{1},\dots,\mu_{m}\in\mathbb{S}^{d}$ and shape parameters $\kappa_{1},\dots,\kappa_{m}>0$, we can define the vMF mixture density with $m$ components by
$f(x)=\sum_{j=1}^{m}w_{j}\cdot c_{d,\kappa_{j}}\exp{(\kappa_{j}\langle x,\mu_{j}\rangle)},\quad x\in\mathbb{S}^{d},\quad d=1,2,$ (5.4)
where $c_{d,\kappa_{j}}$ is the normalisation constant as before. This gives us a flexible family of distributions on the $d$-sphere $\mathbb{S}^{d}$. Let us denote by $W=(w_{1},\dots,w_{m})$, $K=(\kappa_{1},\dots,\kappa_{m})$, and $M=(\mu_{1},\dots,\mu_{m})$, the vectors of weights, shape parameters, and mean directions, respectively.
We plot such a density along with a sample of $n=1000$ points from it and KDEs constructed with $s=0.5$, $s=1$ and $s=2$ in Figure 5. The given density has $m=2$, $W=(\frac{1}{2},\frac{1}{2})$, $K=(12,10)$ and M=((0,0,1),(0,1,0))M=((0,0,1),(0,-1,0)). Table 4 contains the estimated probabilities for regions surrounding each of the two peaks, and corresponding estimated MISE.
Figure 5: (Left to right) The plot of a mixture density over $\mathbb{S}^{2}$; a sample of $n=1000$ points from such a distribution; KDEs constructed using said data for $s=0.5$, $s=1$ and $s=2$, respectively.
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
$N_{s}=19$ $N_{s}=8$ $N_{s}=4$
$0,\pi/6]\times(-\pi,\pi]$ 0.3931 0.368 0.2324 0.394 0.4001
$0,\pi/4]\times(-\pi,\pi]$ 0.4837 0.5024 0.4271 0.482 0.4886
$\pi/3,2\pi/3]\times[-2\pi/3,-\pi/3]$ 0.4022 0.3854 0.2654 0.396 0.4012
$\pi/4,3\pi/4]\times[-3\pi/4,-\pi/4]$ 0.4925 0.5021 0.466 0.492 0.4869
MISE 0.0058 0.017593 0.15324 - -
Table 4: Estimated probabilities from a vMF mixture density on $\mathbb{S}^{2}$ for $n=1000$ points and $s=0.5$, $s=1$, $s=2$.
Unlike the last two examples, both the MISE and plot show that $s=2$ is too large to provide a good fit in this case. This is expected as the density has more local behaviour since it is a mixture density and each component vMF density has a higher shape parameter than before. Indeed, the $s=1$ plot and MISE indicate a better performance than $s=2$, with $s=0.5$ appearing as the best of the three. Moreover, the $s=2$ probabilities are much less accurate than the $s=0.5$ estimations, especially over smaller areas. This is attributed to the oversmoothing we see in this $s=2$ case.
We conclude this section with the estimation of a vMF mixture density over on $\mathbb{S}^{1}$. In this density we take $m=4$ components with weights $W=(\frac{1}{5},\frac{3}{10},\frac{1}{10},\frac{2}{5})$, shape parameters $K=(4,6,10,12)$, and mean directions $M=((1,0),(\frac{1}{2},\frac{\sqrt{3}}{2}),(\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}),(0,-1))$. The corresponding plots and sample are given in Figure 6, with probabilities and MISE given in Table 5. Whilst the estimated probabilities are comparable for each value of $s$, the MISE and plots indicate that $s=1$ provides the best fit of the three. The $s=2$ plot shows oversmoothing and the $s=0.5$ plot shows undersmoothing.
Figure 6: (Left to right) The plot of a mixture density over $\mathbb{S}^{1}$; a sample of $n=1000$ points from such a distribution; KDEs constructed using said data for $s=0.5$, $s=1$ and $s=2$, respectively.
Region $s=0.5$ $s=1$ $s=2$ Frequency True Prob.
$N_{s}=85$ $N_{s}=17$ $N_{s}=6$
$\pi/4,\pi/4]$ 0.717 0.7213 0.7227 0.717 0.6968
$3\pi/4,-\pi/4]$ 0.5316 0.5285 0.5178 0.532 0.5394
$\pi/2,0]$ 0.6057 0.6091 0.615 0.602 0.5999
$\pi/12,7\pi/12]$ 0.6469 0.647 0.6448 0.646 0.659
MISE 0.00824 0.00253 0.01076 - -
Table 5: Estimated probabilities from a vMF mixture density on $\mathbb{S}^{1}$ for $n=1000$ points and $s=0.5$, $s=1$, $s=2$.
5.3 Comparison of Theorem 4.1 with numerical integration
Let us return to the motivational path to Theorem 4.1 and do the next simple experiment.
Consider $n=1000,2000,\dots,10000$ data simulated uniformly on the sphere $\mathbb{S}^{2}$. The finite estimator (3.4) was considered with a symbol $g_{r}(\lambda)$, as in Remark 3.1, where we took $s=1$ and $r=6$. The region of integration was $\pi/4,3\pi/4]\times[-\pi/2,\pi/2]$. We can see in Figure 7 that the closed form, established in Theorem 4.1, significantly outperforms, in terms of time execution, the numerical integration method. For this simulation we used Mathematica’s GlobalAdaptive method with the GaussKronrodRule, the default strategy of the native NIntegrate function.
Figure 7: Evaluation times of the integral of the finite KDEs over a region $\mathbb{S}^{2}$, using numerical integration (red) versus a closed form (blue) for $n\in\{1000,\dots,10\,000\}$.
We summarize the findings of the Section. We have demonstrated the efficacy of our finite KDEs and probability estimation methods in several examples. The outcomes from these simulations were all in line with the theoretical results presented earlier. Moreover, as we noticed, small differences in the values of the selected $s$ give very similar results.
Very large values of $s$, is better to be avoided in a data analysis, because tend to oversmooth.
Very local phenomena are better expressed with values of $s<1$.
With theoretical and empirical evidence for the accuracy of these estimators in hand, we now turn to several case studies with real-world datasets in order to demonstrate how these methods can be used in several research fields.
6 Case Studies
In this section we present four case studies, two on the unit circle $\mathbb{S}^{1}$, and two on the unit sphere $\mathbb{S}^{2}$. We illustrate how kernel density estimators on $\mathbb{S}^{1}$ are well suited for spatially and temporally periodic data. In particular, we analyze the orientations of honeybees and the annual precipitation cycle in Los Angeles. For data on $\mathbb{S}^{2}$, we construct KDEs for global datasets: earthquake locations and the directional distribution of stars in our galaxy. For each dataset, we compute a suitable finite–order KDE, as described in Theorem 3.1, and estimate probabilities over regions of interest using Proposition 4.2 for $\mathbb{S}^{1}$ and Theorem 4.1 for $\mathbb{S}^{2}$.
6.1 Spatially Periodic Data: Honeybee Orientations
A natural source of periodic data is a random variable representing an angle or direction in physical space. Examples include wind and ocean current directions, rock core orientations, and animal movement headings. We illustrate this with the “Honeybee Tracking I” dataset, made available by the Okinawa Institute of Science and Technology https://groups.oist.jp/bptu/honeybee-tracking-dataset. In this dataset, honeybees were filmed at 30 frames per second, and the orientation angle of each bee was recorded at every frame. Here, we analyze the first frame of “Recording 1”, originally gathered for the study [8]. Furthermore, the study [9] also gathered the trajectories of honeybees in a hive with a similar dataset available, also at the above URL.
In this dataset, bees are labelled as 0, if their orientation is clear, or $1$, if their body is partially obscured within the hive. For consistency in our study, we only look at the $n=691$ honeybee orientations of class 0, which are given in degrees. Converting these to radians makes these observations suitable for a KDE as in Theorem 3.1 and Remark 3.1. We selected $s=1$, and therefore $h=n^{-1/(2s+1)}\approx 0.113$, $N_{s}=14$ and $r=5$. The data and corresponding KDE are plotted in Figure 8. In this frame, the bees exhibit a clear preference for facing approximately towards the angle 0.09-0.09 radians with the least favoured p
Figure 8: Plots of $n=691$ honeybee orientations (left) and the corresponding KDE with $s=1$ (right).
In Table 6 we give the estimated probabilities over two regions. The first is [1.33,0.81][-1.33,0.81], which we estimate, from the KDE, to be the main ‘peak’ of the density. The second region is $2.46,\pi]\cup(-\pi,-3.03]$, which we estimate to be the ‘valley’, or the region surrounding the global minimum. Note that for $2.46,\pi]\cup(-\pi,-3.03]$, the estimated probability over this region is given by $\widehat{\mathbb{P}}(A[2.36,\pi])+\widehat{\mathbb{P}}(A(-\pi,-3.03])$, where $A[\vartheta_{1},\vartheta_{2}]$ is given in (4.2). We can see that the empirical and KDE estimated probabilities are approximately equal around the minimum and are relatively close around the maximum.
Region Frequency KDE Prob.
[1.33,0.81][-1.33,0.81] $0.4906$ $0.4893$
$2.46,\pi]\cup(-\pi,-3.03]$ $0.0637$ $0.0637$
Table 6: Empirical and KDE estimated probabilities of honeybee orientations around the maximum and minimum.
Kernel density estimators on $\mathbb{S}^{1}$ were particularly necessary in this case, rather than KDEs over $\mathbb{R}$ as the periodicity of the data needed to be captured, both to properly represent the nature of directional data but also because in this case, the maximum occurred almost at 0.
6.2 Temporally Periodic Data: Annual Precipitation Patterns
We now turn to periodic precipitation patterns at a fixed location, illustrating our methodology for treating temporal phenomena as circular data. The Open-Meteo API provides historical weather records for many locations worldwide; https://open-meteo.com/en/docs/historical-weather-api.
For Los Angeles, California, precipitation follows a clear annual cycle, with a wet season in spring and dry summers. From Open-Meteo, we extracted the days with non-zero precipitation from 01.01.2000 until 01.01.2025 (9133 observed days). This gave $n=1455$ days of precipitation in the time frame. Because the wet season occurs at approximately the same time each year, it is natural to view the data as periodic.
More generally, if a random variable $T$ is periodic on $(a,b]$ with density $f_{T}:(a,b]\to[0,\infty)$ satisfying $f_{T}(a)=f_{T}(b)$, then independent observations $T_{1},\dots,T_{n}\in(a,b]$ may be mapped to circular data on $\mathbb{S}^{1}$ by
$(a,b]$ $\to(-\pi,\pi],$
$t$ $\mapsto\frac{2\pi}{b-a}(t-a)-\pi.$
This transformation allows periodic temporal measurements to be analyzed using kernel density estimation on $\mathbb{S}^{1}$, in line with our theory.
In our case, we assume that rainfall is periodic and therefore that the probability of rain occurring a second after 23:59:59 on the $31^{\text{st}}$ of December is the same as at 00:00:00 on the $1^{\text{st}}$ of January. We then can convert our observations of days of the year with rain occurring into angles on the circle $\mathbb{S}^{1}$, via the above. Note that we adjust the period to be $366$ days in length on leap years.
With this converted data $\theta_{1},\dots,\theta_{1455}$ in hand, we construct a KDE as in (3.3), taking $s=2$, $h\approx 0.233$, $N_{s}=6$ as in Theorem 3.1, and $r=6$. The data, alongside such a KDE is plotted in Figure 9 with the first day of each month indicated on the plot. We can see that the wet spring season is clearly highlighted by the KDE as well as the drier summers with less rain.
We applied the KDE from (3.3) to the angular rainfall data $\theta_{1},\dots,\theta_{1455}$, implementing the finite sum approximation from Theorem 3.1 with parameters $s=2$, $h\approx 0.233$, $N_{s}=6$, and $r=6$. The resulting probability density, plotted in Figure 9 alongside the raw circular data, identifies the climatological regime of Los Angeles. A pronounced mode in the KDE during the spring accurately reflects the core of the wet season, while a broad minimum throughout the summer months captures the extended dry season. The first day of each month is marked on the plot, providing a calendar reference for this seasonal phenomenon.
Figure 9: Plots of $n=1455$ days with nonzero precipitation in Los Angeles (left) and the corresponding KDE with $s=2$ (right).
In Table 7 we give the estimated probabilities of rainfall over the period from the start of February to the end of March (the wet season), as well as the start of June to the end of October (the dry season). These KDE estimated probabilities are compared to the empirical probabilities in those regions. Both methods of probability estimation seem to agree in this table. We also performed this case study for $s=1$, finding that the estimated probabilities were approximately equal as to the $s=2$ case, with the difference being on the order of $10^{-3}$.
Dates (Inclusive) Frequency KDE Prob.
$1^{\text{st}}$ Feb. - $31^{\text{st}}$ Mar. $0.2838$ $0.2815$
$1^{\text{st}}$ Jun. - $31^{\text{st}}$ Oct. $0.2027$ $0.2008$
Table 7: Empirical and KDE estimated probabilities of rainfall in LA over certain periods.
By viewing seasonally dependent events as occurring on $\mathbb{S}^{1}$, we are able to capture their periodic nature. If instead we were to view this data as having occurred over $\mathbb{R}$ there would be two problems. Firstly, the kernel density estimator would estimate a very low probability for a date in 1990 or 2030, since it has no data for either period. Secondly, a particularly dry or wet year, i.e. an outlier, would greatly skew the estimation over that period of time, since there is at most one observation at any given day of the year. One could attempt to tackle both of these problems by using a KDE on a finite, non-periodic interval, say $[0,1]$ and scaling each observation accordingly. However, a KDE on this metric space wouldn’t be periodic and there would likely be a discontinuity in the estimator as it moves from the 31st of December to the 1st of January. Therefore, the use of kernel density estimators over $\mathbb{S}^{1}$ is essential in order to address the above problems and to properly construct an estimator of such a phenomenon.
6.3 Global Data: Earthquake Locations
An obvious setting corresponding to the sphere $\mathbb{S}^{2}$ is the surface of the Earth. With this in mind, we present an application of kernel density estimation to events on $\mathbb{S}^{2}$ in estimating the distribution of earthquakes, as discussed in [14]. Global earthquake data is freely available through the United States Geological Survey website https://earthquake.usgs.gov/earthquakes/search/.
There were $n=1630$ earthquakes of magnitude $6.5$ or higher from 1990 to 2024 (inclusive). Each earthquake recording comes with its latitude and longitude, which can easily be converted into Cartesian coordinates on $\mathbb{S}^{2}$. Indeed, for a given latitude (lat) and longitude (lon) one can find the corresponding point on the sphere $X\in\mathbb{S}^{2}$ via the following:
$X=\Big(\cos\Big(\frac{\text{lat}\cdot\pi}{180}\Big)\cos\Big(\frac{\text{lon}\cdot\pi}{180}\Big),\cos\Big(\frac{\text{lat}\cdot\pi}{180}\Big)\sin\Big(\frac{\text{lon}\cdot\pi}{180}\Big),\sin\Big(\frac{\text{lat}\cdot\pi}{180}\Big)\Big).$
With the above conversion in hand, we have observations $X_{1},\dots,X_{1603}\in\mathbb{S}^{2}$. We select $s=0.05$ and construct a KDE $\widehat{f}_{s}$, as in Theorem 3.1 and Remark 3.1, taking $h=n^{-1/(2s+2)}\approx 0.0295$, $N_{s}=92$ and $r=6$. The corresponding KDE as well as the data is plotted in Figure 10.
Figure 10: Plots of $n=1630$ high magnitude earthquakes (left) and the corresponding KDE with $s=0.05$ (right).
The resulting KDE effectively delineates regions of high seismic activity, with the Pacific ’Ring of Fire’ appearing as the most prominent global feature. The model successfully captures the complex tectonics of Indonesia, where the Ring of Fire intersects with the Alpide Belt. We can use the probability estimation described in Section 4 to estimate the probabilities over regions of particularly high (and low) seismic activity. We computed the probability of a high-magnitude earthquake within the bounding boxes of Japan, Chile, the Philippines, and Ireland by integrating the KDE over these regions. The results, presented in Table 8 alongside the corresponding bounding coordinates and observed frequencies (to four decimal places), provide a probabilistic characterization of seismic hazard in these locations.
Note that when evaluating (4.6) for high cutoff points (about $N_{s}\geq 60$, depending on the machine precision), it is recommended to using exact numbers. The main reason for this is due to the numerical ill-conditioning of high-degree orthogonal polynomials, due to catastrophic cancellation. We found that using exact numbers was the simplest method to avoid this issue, however, there also exists numerically stable algorithms as alternatives. As an example, Mathematica distinguishes between the Real (floating-point) number 0.9 and the Rational number 9/10. The evaluation of an expansion of LegendreP[60,x] gives -512 at x=0.9, clearly incorrect, whereas for x=9/10 it gives 0.0317896.
Country Min. Lat. Max. Lat. Min. Lon. Max. Lon. Frequency KDE Prob.
Japan $31^{\circ}$N $45.5^{\circ}$N $129.4^{\circ}$E $145.5^{\circ}$E $0.0540$ $0.049808$
Chile $55.6^{\circ}$S $17.6^{\circ}$S $75.6^{\circ}$W $70^{\circ}$W $0.0429$ $0.043289$
Philippines $5.6^{\circ}$N $18.5^{\circ}$N $117.2^{\circ}$E $126.5^{\circ}$E $0.0270$ $0.025305$
Ireland $51.7^{\circ}$N $55.1^{\circ}$N $10^{\circ}$W $6^{\circ}$W $0.0000$ $0.000002$
Table 8: Empirical and KDE estimated probabilities of high magnitude earthquakes occurring within the bounding boxes of four regions on Earth.
This application highlight’s the importance of having kernel density estimators available on $\mathbb{S}^{2}$. If one were to have instead constructed an estimator over $\mathbb{R}^{2}$, using the latitude and longitude as coordinates, the periodicity in the longitude would be lost. For instance, there would be a discontinuity between the earthquake density from Eastern Russia to Alaska. Furthermore, the distances between, say $(90^{\circ}\text{N},20^{\circ}\text{E})$ and $(90^{\circ}\text{N},30^{\circ}\text{E})$ would be taken by the KDE to be the same as the distance between $(0^{\circ}\text{N},20^{\circ}\text{E})$ and $(0^{\circ}\text{N},30^{\circ}\text{E})$, when in fact the former two points coincide. This would cause an underestimation of the density near the poles.
6.4 Astronomical Data: Star Directions
Distributions occurring on the sphere $\mathbb{S}^{2}$ are not limited to just phenomena on Earth. Indeed, to each visible star, astronomers assign a “galactic longitude” and “galactic latitude”. This galactic coordinate system assigns the sun of our solar system as its origin. The galactic equator is traced out from the distance of our sun to the approximate centre of the Milky Way. Using the right-hand convention, galactic longitude and galactic latitude are analogues of the familiar terrestrial coordinates and can therefore be converted to Cartesian coordinates via (6.3).
The Yale Bright Star Catalogue, freely available at http://tdc-www.harvard.edu/catalogs/bsc5.html, is a commonly used dataset of stars brighter than magnitude $6.5$. It contains the galactic longitude and latitude of $n=9096$ bright stars. As before, these can be easily converted into Cartesian coordinates $X_{1},\dots,X_{9096}\in\mathbb{S}^{2}$.
Figure 11 presents the Bright Star Catalogue dataset alongside its (KDE) computed with $s=1$. This specific $s$ value was chosen to suppress small-scale noise while still resolving the major structural features of the distribution. The resulting KDE map, rendered in the standard galactic coordinate system where Earth is located at $(0^{\circ},0^{\circ})$ latitude and longitude, reveals regions of increased stellar density. In accordance with astronomical convention, the galactic longitude axis is reversed in the plots. A prominent band of high density is observed along the mid-plane of the Galaxy, an expected result that reflects the concentration of bright stars within the galactic disc. This concentration arises both from the higher density of stars in the disc and from the presence of luminous, massive stars which are largely confined to this plane.
Figure 11: Plots of $n=9096$ bright stars (left) and the corresponding KDE with $s=1$ (right).
In Table 9 we use Theorem 4.1 in order to estimate the probabilities over certain regions of space using the KDE. Indeed, we present the probabilities over the four quadrants of the galaxy, as given by galactic coordinates, as well as the probability over what appears to be the mode of the distribution, which we assume has a bounding box given by $(20^{\circ}$S, $130^{\circ}$W )) and $(5^{\circ}$N, $80^{\circ}$ W)). In the table we also present the empirical probabilities in each region. One can see that both the empirical probability and the KDE probability closely align.
Region Min. Lat Max. Lat. Min. Lon. Max. Lon. Frequency KDE Prob.
Peak $20^{\circ}$S $5^{\circ}$N $80^{\circ}$W $130^{\circ}$W $0.0629$ $0.0607$
Upper Left $0^{\circ}$N $90^{\circ}$N $180^{\circ}$W $0^{\circ}$E $0.233$ $0.2368$
Upper Right $0^{\circ}$N $90^{\circ}$N $0^{\circ}$E $180^{\circ}$E $0.2379$ $0.2407$
Lower Left $90^{\circ}$S $0^{\circ}$N $180^{\circ}$W $0^{\circ}$E $0.2898$ $0.2847$
Lower Right $90^{\circ}$S $0^{\circ}$N $0^{\circ}$E $180^{\circ}$E $0.2392$ $0.2379$
Table 9: The estimated (empirical and KDE) probabilities over certain regions of the sphere as well as the bounding coordinates of the regions, given in spherical coordinates.
This case study demonstrates the usefulness of kernel density estimators over $\mathbb{S}^{2}$ to estimate the distribution of the directions of vectors in $\mathbb{R}^{3}$, where magnitude is irrelevant.
7 Conclusion
In our study we provided nonparametric density and probability estimators on the circle $\mathbb{S}^{1}$ and the sphere $\mathbb{S}^{2}$ that are both rate-optimal and computationally efficient. The key idea is to replace the usual infinite expansions underlying kernel density estimators with their finite-order counterparts, determining the order index in accordance with the smoothness of the density.
Moreover, our probability estimators admit closed-form expressions, which enables rapid computation even for large datasets. This strengthens the applicability of kernel methods in the analysis of circular and spherical data, where region-based probability summaries are often of primary interest.
We demonstrated a general methodology applicable to periodic and directional datasets and applied it in problems arising from zoology, climate science, seismology, and astronomy. These examples highlight the interpretability and computational advantages of working directly in the natural geometry of the data, rather than projecting it to Euclidean space. Our methodologies can be readily applied in a broad range of sciences.
There are several natural directions for future research. One is the extension of the finite-order methodology to higher-dimensional spheres $\mathbb{S}^{d}$, $d>2$, or other metric spaces associated with Laplacians with discrete spectrum as in [10]. Another very natural direction is the study of adaptive adaptive estimators that do not rely on prior smoothness assumptions, something that is in our future agenda. Finally, Bayesian nonparametric models for directional data may provide a promising route for future research.
Acknowledgements
The first-named author would like to thank Prof. Morgan Fraser for his helpful comments and suggestions on the astronomical application discussed in Section 6.4. The second-named author would like to thank Prof. Dr. Katarzyna Bozek an author of [8] and [9], whose dataset and expertise enabled the discussion in zoology in Section 6.1.
8 Appendix
This section contains the proofs of our results.
8.1 Proof of Theorem 3.1
We start by proving the optimal rate of the finite estimators introduced in this study.
Proof.
The proof presents similarities, but also some differences for the case $d=1$ and $d=2$. Thus, we distinguish these cases.
Case $d=2$.
Consider the MSE of our new estimators in (3.4) and use the trivial inequality $(a+b)^{2}\leq 2(a^{2}+b^{2})$;
$\text{MSE}(\widehat{f}_{s})(x)$ $\leq 2\left(\mathbb{E}\Big[\big(f(x)-\widehat{f}_{n,h}(x)\big)^{2}\Big]+\mathbb{E}\Big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\Big]\right)$ (8.1)
$=2\left(\text{MSE}(\widehat{f}_{n,h})(x)+\mathbb{E}\Big[\big(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big)^{2}\Big]\right).$
[…]
$\big|\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\big|=\Big|\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell>N_{s}}\frac{2\ell+1}{4\pi}g(h\sqrt{\ell(\ell+1)})P_{\ell}(\langle x,X_{j}\rangle)\Big|.$ (8.2)
By (3.1) we take
$|g(h\sqrt{\ell(\ell+1)})|\leq C_{\tau}(1+h\sqrt{\ell(\ell+1)})^{-r}<C_{\tau}h^{-r}\big(\ell(\ell+1)\big)^{-r/2}.$
For the Legendre polynomials we have $|P_{\ell}(u)|\leq 1,$ for every $u\in[-1,1]$; see (2.3).
Then the error under consideration can be bounded from above by
$|\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)|$ $\leq\frac{C_{\tau}}{4\pi}h^{-r}\sum_{\ell=N_{s}+1}^{\infty}(2\ell+1)\big(\ell(\ell+1)\big)^{-r/2}$
$\leq\frac{C_{\tau}}{4\pi}h^{-r}\int_{N_{s}}^{\infty}(2x+1)\big(x(x+1)\big)^{-r/2}dx$
$=\frac{C_{\tau}}{2\pi(r-2)}h^{-r}\big(N_{s}(N_{s}+1)\big)^{1-\frac{r}{2}}$
$\leq\frac{C_{\tau}}{2\pi(r-2)}h^{-r}N_{s}^{2-r}.$ (8.3)
Note that above we used the usual bound of the series of decreasing sequences by the corresponding integral and then we changed the variable of integration.
For the choice of the bandwidth it holds $h^{-r}=n^{r/(2s+2)}$.
By the selected $N_{s}$ as in (3.2) we get
$N_{s}^{2-r}\leq 2\pi(r-2)n^{-(s+r)/(2s+2)}.$
Replacing in (8.1) we derive the bound
$|\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)|\leq C_{\tau}n^{-s/(2s+2)}.$ (8.4)
Combining (8.1) with (8.4) and the result from Theorem 2.1 we arrive at (3.5) and complete the proof for $d=2$.
Case $d=1$. As in the previous case we focus in $|\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)|$.
For a given $\ell\in\mathbb{N}$, we have using (3.1) that
$|g(h\ell)|\leq C_{\tau}(1+h\ell)^{-r}<(h\ell)^{-r}.$ (8.5)
Moreover, $|\cos(t)|\leq 1$ for all $t\in\mathbb{R}$. Therefore, we have the following bound:
$|\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)|$ $\leq\frac{1}{\pi n}\sum_{j=1}^{n}\sum_{\ell=N_{s}+1}^{\infty}h^{-r}\ell^{-r}=\frac{h^{-r}}{\pi}\sum_{\ell=N_{s}+1}^{\infty}\ell^{-r}$
$\leq\frac{h^{-r}}{\pi}\int_{N_{s}}^{\infty}u^{-r}du=\frac{h^{-r}}{\pi}\frac{N_{s}^{1-r}}{r-1},$
for every $x\in\mathbb{S}^{1}$.
Now consider $N_{s}$ as in (3.2). One has that
$N_{s}^{1-r}\leq\pi(r-1)n^{-(s+r)/(2s+1)},$
and substituting $h=n^{-1/(2s+1)}$, we get
$|\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)|\leq n^{-s/(2s+1)}.$
and therefore, we have that
$\mathbb{E}\left[\left(\widehat{f}_{n,h}(x)-\widehat{f}_{s}(x)\right)^{2}\right]\leq n^{-2s/(2s+1)}.$
Combining this with (8.1) and our bound for the MSE of $\widehat{f}_{n,h}$ from Theorem 2.1 gives the result for $d=1$ and completes the proof. ∎
8.2 Proofs of the results in probability estimation
We present here all proofs of the results in Section 4. We start as follows:
Proof.
(of Proposition 4.1). The probability of the measurable set $A$ is
$\mathbb{P}(A)=\int_{A}f(x)d\sigma_{\mathbb{S}^{d}}(x).$
Then by the Cauchy-Schwarz inequality for $\mathcal{L}^{2}(\mathbb{S}^{d})$ and Fubini-Tonelli Theorem we have that
$\mathbb{E}_{f}$ $\big(\widehat{\mathbb{P}}(A)-\mathbb{P}(A)\big)^{2}\leq\mathbb{E}_{f}\Big(\int_{A}\big|\widehat{f}_{s}(x)-f(x)\big|d\sigma_{\mathbb{S}^{d}}(x)\Big)^{2}$
$\leq\sigma_{\mathbb{S}^{d}}(A)\mathbb{E}_{f}\int_{A}\big(\widehat{f}_{s}(x)-f(x)\big)^{2}d\sigma_{\mathbb{S}^{d}}(x)$
$=\sigma_{\mathbb{S}^{d}}(A)\int_{A}\mathbb{E}_{f}\big(\widehat{f}_{s}(x)-f(x)\big)^{2}d\sigma_{\mathbb{S}^{d}}(x)\leq\sigma_{\mathbb{S}^{d}}(A)^{2}\sup_{x\in\mathbb{S}^{d}}\text{MSE}(\widehat{f}_{s}(x)),$
and the result follows from Theorem 3.1.
The main probability estimation on $\mathbb{S}^{1}$; Proposition 4.2 has the following proof:
Proof.
We are concerned with evaluating $\int_{A[\vartheta_{1},\vartheta_{2}]}\widehat{f}(x)d\sigma_{\mathbb{S}^{1}}(x)=\int_{\vartheta_{1}}^{\vartheta_{2}}\widehat{f}_{s}(\theta)d\theta$ where $\widehat{f}_{s}$ is as in (3.3). By simple integration we confirm the statement:
$\widehat{\mathbb{P}}$ $(A[\vartheta_{1},\vartheta_{2}])=\int_{\vartheta_{1}}^{\vartheta_{2}}\frac{1}{2\pi n}\sum_{j=1}^{n}\left(1+2\sum_{\ell=1}^{N_{s}}g(h\ell)\cos(\ell(\theta-\theta_{j}))\right)d\theta$
$=\frac{\vartheta_{2}-\vartheta_{1}}{2\pi}+\frac{1}{\pi n}\sum_{j=1}^{n}\sum_{\ell=1}^{N_{s}}g(h\ell)\int_{\vartheta_{1}}^{\vartheta_{2}}\cos(\ell(\theta-\theta_{j}))d\theta.$
$=\frac{\vartheta_{2}-\vartheta_{1}}{2\pi}+\frac{1}{\pi n}\sum_{j=1}^{n}\sum_{\ell=1}^{N_{s}}g(h\ell)\frac{1}{\ell}(\sin(\ell(\vartheta_{2}-\theta_{j}))-\sin(\ell(\vartheta_{1}-\theta_{j}))).$
Let us now proceed to prove the Lemma 4.2 on the integrals of the associated Legendre functions.
Proof.
Recall
First we combine (2.4) and (2.2). For every $\ell\in\mathbb{N}$ and $1\leq m\leq\ell$, we differentiate (2.2) to extract
$\frac{d^{m}}{du^{m}}P_{\ell}(u)$ $=\sum_{k=m}^{\ell}{\scriptstyle\binom{\ell}{k}\binom{\ell+k}{k}}\frac{(-1)^{k}}{2^{k}}\frac{d^{m}}{du^{m}}(1-u)^{k}$
$=\sum_{k=m}^{\ell}{\scriptstyle\binom{\ell}{k}\binom{\ell+k}{k}}\frac{(-1)^{k+m}}{2^{k}}k(k-1)\cdots(k-m+1)(1-u)^{k-m}$
$=\sum_{k=m}^{\ell}{\scriptstyle\binom{\ell}{k}\binom{\ell+k}{k}}\frac{(-1)^{k+m}}{2^{k}}\frac{k!}{(k-m)!}(1-u)^{k-m}.$ (8.6)
We embed it into (2.4) and obtain for every $\ell\geq 1,\;1\leq m\leq\ell$
$P_{\ell}^{m}(u)$ $=(-1)^{m}(1-u^{2})^{m/2}\frac{d^{m}}{du^{m}}P_{\ell}(u)$ (8.7)
$=\sum_{k=m}^{\ell}{\scriptstyle\binom{\ell}{k}\binom{\ell+k}{k}}\frac{(-1)^{k}}{2^{k}}\frac{k!}{(k-m)!}(1+u)^{m/2}(1-u)^{k-m/2}.$
Therefore for the integral $\int_{\gamma_{1}}^{\gamma_{2}}P_{\ell}^{m}(u)du$, it suffices to compute $\int_{\gamma_{1}}^{\gamma_{2}}(1+u)^{m/2}(1-u)^{k-m/2}du$. We consider the more general integral
$O(a,b,\gamma_{1},\gamma_{2}):=\int_{\gamma_{1}}^{\gamma_{2}}(1+u)^{a}(1-u)^{b}du,\quad a,b>-1.$ (8.8)
We change to the variable $t:=\frac{1+u}{2}$ and for $1\leq\gamma_{1}\leq u\leq\gamma_{2}\leq 1$, we have that $0\leq\frac{1+\gamma_{1}}{2}\leq t\leq\frac{1+\gamma_{2}}{2}\leq 1$ and therefore
$O(a,b,\gamma_{1},\gamma_{2})$ $=2^{a+b+1}\int_{\frac{1+\gamma_{1}}{2}}^{\frac{1+\gamma_{2}}{2}}t^{a}(1-t)^{b}du$
$=2^{a+b+1}\Big[B\Big(\frac{1+\gamma_{2}}{2};a+1,b+1\Big)-B\Big(\frac{1+\gamma_{1}}{2};a+1,b+1\Big)\Big].$ (8.9)
We can now apply (8.9) to (8.7) for $a=m/2$ and b=km/2b=k-m/2 to conclude the proof of the Lemma. ∎
Proof of Theorem 4.1
We can now proceed presenting the proof of the main result in probability estimation.
Proof.
We consider the new optimal estimator in (3.4) and we use the addition formula for Legendre polynomials and spherical harmonics, given in (2.7), taking
$\widehat{f}_{s}(x)$ $=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{N_{s}}\frac{2\ell+1}{4\pi}g\Big(h\sqrt{\ell(\ell+1)}\Big)P_{\ell}(\langle x,X_{j}\rangle)$
$=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{N_{s}}g\Big(h\sqrt{\ell(\ell+1)}\Big)\sum_{m=-\ell}^{\ell}Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})},$ (8.10)
where $(\theta,\phi)$ the spherical coordinates of $x\in\mathbb{S}^{2}$.
Based on the properties of complex numbers and the complex conjugation between spherical harmonics in (2.6) we infer that
$\sum_{m=-\ell}^{\ell}$ $Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}$
$=\sum_{m=-\ell}^{-1}Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}+Y^{0}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{0}(\theta_{j},\phi_{j})}$
$\hskip 28.45274pt+\sum^{\ell}_{m=1}Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}$
$=Y^{0}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{0}(\theta_{j},\phi_{j})}+2\sum^{\ell}_{m=1}\text{Re}\big(Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}\big).$ (8.11)
Observe that for $m=0$, the corresponding spherical harmonics $Y_{\ell}^{0}$ are real-valued and precisely
$Y^{0}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{0}(\theta_{j},\phi_{j})}=N_{\ell,0}^{2}P_{\ell}(\cos\theta)P_{\ell}(\cos\theta_{j}).$ (8.12)
On the other hand, by the formula (2.5) we extract
$\text{Re}\big(Y^{m}_{\ell}(\theta,\phi)\overline{Y_{\ell}^{m}(\theta_{j},\phi_{j})}\big)$ $=N_{\ell,m}^{2}P_{\ell}^{m}(\cos\theta)P^{m}_{\ell}(\cos\theta_{j})\text{Re}\Big(e^{im(\phi-\phi_{j})}\Big)$
$=N_{\ell,m}^{2}P_{\ell}^{m}(\cos\theta)P^{m}_{\ell}(\cos\theta_{j})\cos\big(m(\phi-\phi_{j})\big).$ (8.13)
Combining (8.10)-(8.13), the estimated probability takes the form
$\widehat{\mathbb{P}}(\mathcal{A}_{\vartheta,\varphi})$ $=\int_{\mathcal{A}_{\vartheta,\varphi}}\widehat{f}_{s}(x)d\sigma_{\mathbb{S}^{d}}(x)$
$=\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{N_{s}}g\Big(h\sqrt{\ell(\ell+1)}\Big)\times$
$\hskip 28.45274pt\Bigg[\int_{\vartheta_{1}}^{\vartheta_{2}}\int_{\varphi_{1}}^{\varphi_{2}}N_{\ell,0}^{2}P_{\ell}(\cos\theta)P_{\ell}(\cos\theta_{j})\sin\theta d\phi d\theta$
$\hskip 28.45274pt+2\sum_{m=1}^{\ell}\int_{\vartheta_{1}}^{\vartheta_{2}}\int_{\varphi_{1}}^{\varphi_{2}}N_{\ell,m}^{2}P_{\ell}^{m}(\cos\theta)P^{m}_{\ell}(\cos\theta_{j})\times$
$\hskip 113.81102pt\times\cos\big(m(\phi-\phi_{j})\big)\sin\theta d\phi d\theta\Bigg]$
$=:\frac{1}{n}\sum_{j=1}^{n}\sum_{\ell=0}^{N_{s}}g\Big(h\sqrt{\ell(\ell+1)}\Big)\Big(I_{\ell}^{0}(j)+2\sum_{m=1}^{\ell}I_{\ell}^{m}(j)\Big),$ (8.14)
and we must prove that the integrals $I_{\ell}^{m}(j)$, agree with their values in the statement of the theorem.
In the integrals $I_{\ell}^{0}(j)$ first. We observe by changing the variable $u=\cos\theta$ that
$I_{\ell}^{0}(j)=N_{\ell,0}^{2}(\varphi_{2}-\varphi_{1})P_{\ell}(\cos\theta_{j})\int^{\cos\vartheta_{1}}_{\cos\vartheta_{2}}P_{\ell}(u)du.$ (8.15)
Recall the expansion (2.2) of the Legendre polynomials and focus on the integrals $\int^{\cos\vartheta_{1}}_{\cos\vartheta_{2}}(1-u)^{k}du$, $\leq k\leq\ell$.
It is easily observed (we may distinguish the cases $k=0$ and $k>0$ to compute the unified result), that for all $0\leq k\leq\ell$
$\int^{\cos\vartheta_{1}}_{\cos\vartheta_{2}}(1-u)^{k}du=\frac{1}{k+1}\Big((1-\cos\vartheta_{2})^{k+1}-(1-\cos\vartheta_{1})^{k+1}\Big),\quad k\geq 0,$
which concludes the calculation of $I_{\ell}^{0}(j)$ to
$I_{\ell}^{0}(j)$ $=N_{\ell,0}^{2}(\varphi_{2}-\varphi_{1})P_{\ell}(\cos\theta_{j})\sum_{k=0}^{\ell}{\scriptstyle\binom{\ell}{k}\binom{\ell+k}{k}}\frac{(-1)^{k}}{2^{k}}\cdot\frac{1}{k+1}\times$
$\hskip 142.26378pt\left((1-\cos\vartheta_{2})^{k+1}-(1-\cos\vartheta_{1})^{k+1}\right),$
and confirms the corresponding formula in our statement.
Let $\ell\geq 1$ and $1\leq m\leq\ell$ and consider the integral
$I_{\ell}^{m}(j)$ $=\frac{1}{m}N_{\ell,m}^{2}P^{m}_{\ell}(\cos\theta_{j})\Big(\sin\big(m(\varphi_{2}-\phi_{j})\big)-\sin\big(m(\varphi_{1}-\phi_{j})\big)\Big)\times$
$\hskip 142.26378pt\times\int^{\cos\vartheta_{1}}_{\cos\vartheta_{2}}P_{\ell}^{m}(u)du.$ (8.16)
All the attention is finally on the integral
$\int^{\cos\vartheta_{1}}_{\cos\vartheta_{2}}P_{\ell}^{m}(u)du.$
By Lemma 4.2 and (8.2) we obtain
$I^{m}_{\ell}(j)$ $=\frac{1}{m}N_{\ell,m}^{2}P^{m}_{\ell}(\cos\theta_{j})\Big(\sin\big(m(\varphi_{2}-\phi_{j})\big)-\sin\big(m(\varphi_{1}-\phi_{j})\big)\Big)\times$
$\hskip 17.07182pt\sum_{k=m}^{\ell}{\binom{\ell}{k}\binom{\ell+k}{k}}\frac{2(-1)^{k}k!}{(k-m)!}\mathcal{B}_{k,m}(\cos\vartheta_{2},\cos\vartheta_{1}),$
as claimed, and the proof is complete. ∎